OpenStudy (anonymous):

what are the steps to derive (3/4)cos^2(2x)?

OpenStudy (raden):

the chain's rule again :D

OpenStudy (anonymous):

I know, im confused how to use the chain rule with cos^2, though :/

OpenStudy (raden):

(2x) ' = 2 ((cos)^2)'= 2 cos (cos)' = -sin then times them

OpenStudy (anonymous):

did not know that the derivative of cos^2 was 2cos lol. that is useful

OpenStudy (raden):

it like derive y^2, get 2y. hehe :)

OpenStudy (raden):

so, what do you get ?

OpenStudy (anonymous):

(3/2)cos2(-sinx)?

OpenStudy (anonymous):

2cos*

OpenStudy (anonymous):

wait idk lol

OpenStudy (anonymous):

(3/2)cosx2(-sinx) maybe?

OpenStudy (raden):

(3/4)cos^2(2x) = (3/4) * 2 * 2cos(2x) * (-sin(2x)) = - 3/4 * 4 * sin(2x) cos(2x) you can cancel the 4's

OpenStudy (raden):

btw, do you have the answer choices ?

OpenStudy (anonymous):

its free response. ok so chain rule is: derivative of the inside * the function * derivative of the function?

OpenStudy (raden):

yes, inside first. the power, then the function

OpenStudy (anonymous):

I have it in my notes as f(g(x)) = f'(g(x))g' does that look right?

OpenStudy (raden):

correct!

OpenStudy (anonymous):

so in my problem f(x) = (cos^2)(x) and g(x) = 2x?

OpenStudy (raden):

yep

OpenStudy (raden):

and actually using any identity of trigonometry to simplify it

OpenStudy (anonymous):

alright I think I understand this problem, thanks again man :D

OpenStudy (raden):

it like i want simplify the answer above : (3/4) * 2 * 2cos(2x) * (-sin(2x)) = - (3/2) * 2cos(2x) * sin(2x) = -3/2 sin(4x) you're welcome friend :)